OpenStudy (anonymous):

Find the sum of three consecutive odd integers if the sum of the first two integers is equal to twenty-four less than four times the third integer.

OpenStudy (anonymous):

2n + 2 = 4(n+4) - 24

OpenStudy (anonymous):

sorry half my answer deleted itself there

OpenStudy (anonymous):

let the first odd number = n the sum, s, is given by: n + n+2 + n+4 = s the last bit of information tells us n + n+2 = 4(n+4) - 24

OpenStudy (anonymous):

solve the last equation to find n, then use this value to find s

OpenStudy (anonymous):

Thank you SO much! :D!

OpenStudy (anonymous):

no problem :)